OpenStudy (anonymous):

how to integrate 4cos^2(3x)

ganeshie8 (ganeshie8):

rewrite \(\cos^2x\) as \((\cos(2x) + 1) / 2\)

OpenStudy (shamim):

u hv to use a trigonometric formula here 2cos^A=1+cos2A

OpenStudy (anonymous):

how to integrate 4cos6x+2 then?

ganeshie8 (ganeshie8):

\(\large \int 4\cos^2(3x) ~dx\) \(\large \int 2\times 2\cos^2(3x) ~dx\) \(\large \int 2( 1 + \cos (2\times 3x)) ~dx\) \(\large \int 2 + 2\cos (6x) ~dx\)

ganeshie8 (ganeshie8):

you got that right ?

OpenStudy (anonymous):

Yeah I got 2+2cos6x. sorry for the mistake earlier ><

ganeshie8 (ganeshie8):

\(\large \int 2 + 2\cos (6x) ~dx\) split the integral : \(\large \int 2~ dx + \int 2\cos (6x) ~dx\) \(\large 2\int 1~ dx + 2\int \cos (6x) ~dx\)

ganeshie8 (ganeshie8):

whats the integral : \(\int 1~ dx\) ?

OpenStudy (anonymous):

x?

ganeshie8 (ganeshie8):

yes, wat about the integral : \(\int \cos (x) ~dx\) ?

OpenStudy (anonymous):

sin x?

ganeshie8 (ganeshie8):

yes, but since we have cos(3x), we need to fix it a bit : \(\int \cos(3x)~dx = \frac{\sin(3x)}{3} + c\)

ganeshie8 (ganeshie8):

So, \(\large 2\int 1~ dx + 2\int \cos (6x) ~dx\) evaluates to : \(\large 2x + 2\frac{\sin(6x)}{6} + C\)

ganeshie8 (ganeshie8):

simplify if u can

OpenStudy (anonymous):

Omg, thank you so much!!!

ganeshie8 (ganeshie8):

u wlc :)